Arrive at the Lewis structure of PCl5 and XeF4 using the steps indicated in
your Unit.
1
Expert's answer
2019-02-07T03:11:44-0500

1. PCl5

Centarl atom is P

Number of valence electrons of P is 5

Number of valence electrons of Cl is 7

Total number of valence electrons in the molecule is 5 + 7*5 = 40 valence electrons

Place the bonding pair of electrons between each pair of adjacent atoms to give a single bond. Beginning with the terminal atoms, add enough electrons to each atom to give each atom an octet (we will get Lewis structure of PCl5):

Place the bonding pair of electrons between each pair of adjacent atoms to give a single bond. Beginning with the terminal atoms, add enough electrons to each atom to give each atom an octet (we will get Lewis structure of PCl5):

Formal Charge = valence electrons (free atom) - (nonbonding electrons + bonding electrons/2)

Formal charge (P) = 5 - (0 + 10/2) = 0

Formal Charge (Cl) = 7 - (6 + 2/2) = 0

Formal charges on all the atoms = 0 , this structure is the most stable.

2. XeF4

Centarl atom is Xe

Number of valence electrons of Xe is 8

Number of valence electrons of F is 7

Total number of valence electrons in the molecule is 8 + 7*4 = 36 valence electrons

Place the bonding pair of electrons between each pair of adjacent atoms to give a single bond. Beginning with the terminal atoms, add enough electrons to each atom to give each atom an octet, if any electrons are left over, place them on the centarl atom. Centarl atom is Xe (it it in the 5th row) can have an expanded octet (we will get Lewis structure of XeF4):

Place the bonding pair of electrons between each pair of adjacent atoms to give a single bond. Beginning with the terminal atoms, add enough electrons to each atom to give each atom an octet, if any electrons are left over, place them on the centarl atom. Centarl atom is Xe (it it in the 5th row) can have an expanded octet (we will get Lewis structure of XeF4):

Formal Charge = valence electrons (free atom) - (nonbonding electrons + bonding electrons/2)

Formal charge (Xe) = 8 - (4 + 8/2) = 0

Formal Charge (F) = 7 - (6 + 2/2) = 0

Formal charges on all the atoms = 0 , this structure is the most stable.